OpenStudy (anonymous):

Find the domain for 2x^2+7x-5, which is the denominator for a certain fraction.... I did the quadratic formula, and got a weird answer.... can someone please help me?

jimthompson5910 (jim_thompson5910):

what did you get when you used the quadratic formula?

OpenStudy (anonymous):

hold on, let me find my notes.There's stuff all over my room, I am remodeling, and its somewhere...

OpenStudy (anonymous):

-7+/-sqrt(89)/4

jimthompson5910 (jim_thompson5910):

I'm assuming everything is over 4 If so, then you have the correct answer.

jimthompson5910 (jim_thompson5910):

Those are the zeros. They make the entire polynomial equal to zero.

jimthompson5910 (jim_thompson5910):

So you have to exclude them from the domain

OpenStudy (anonymous):

how would you put that into the form (a,b)?

jimthompson5910 (jim_thompson5910):

You can either enter them as they are or you can approximate them So if you use the exact answers, then the domain is \[ \left(-\infty, \frac{-7-\sqrt{89}}{4}\right)\cup\left(\frac{-7-\sqrt{89}}{4}, \frac{-7+\sqrt{89}}{4}\right)\cup\left(\frac{-7+\sqrt{89}}{4}, \infty\right)\] but that is really really messy.... so I'd go with the approximate zeros The two zeros are approximately -4.108495283 and 0.6084952830 This means we have to exclude these values from the domain So the domain is \[ \left(-\infty, -4.108495283\right)\cup\left(-4.108495283, 0.6084952830\right)\cup\left(0.6084952830, \infty\right)\]

OpenStudy (anonymous):

can you put that in fractional form, please?

jimthompson5910 (jim_thompson5910):

the first answer is in fractional form

jimthompson5910 (jim_thompson5910):

but you can't represent the answer in the form a/b where a and b are integers

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

how about rounding it?

jimthompson5910 (jim_thompson5910):

to the nearest thousandth, the domain is \[ \left(-\infty, -4.108\right)\cup\left(-4.108, 0.608\right)\cup\left(0.608, \infty\right)\]

OpenStudy (anonymous):

I thought the domain for this one was all real numbers.

jimthompson5910 (jim_thompson5910):

not if that polynomial is the denominator of some fraction

OpenStudy (anonymous):

oh.

jimthompson5910 (jim_thompson5910):

yeah that was my initial thought too

myininaya (myininaya):

hopefully the domain of the numerator is all real numbers

jimthompson5910 (jim_thompson5910):

true...

jimthompson5910 (jim_thompson5910):

but I'm assuming that the numerator is some polynomial too